How long will it take to fill a sphere-shaped vat with radius 18.1 ft with a certain liquid (d=2.09 g/mL) if it is spilling out of the hose at 11821.4 g/s?
Answer #1

conversion to be used:

1 ft = 30.48 cm

radius,

r = 18.1 ft

= 18.1 * 30.48 cm

= 551.7 cm

use:

volume = (4/3)*pi*r^3

= (4/3)*3.14*(551.7 cm)^3

= 7.03*10^8 cm^3

= 7.03*10^8 mL

mass of liquid = density * volume

= (2.09 g/mL)*(7.03*10^8 mL)

= 1.47*10^9 g

now use:

time = mass of liquid / rate of spilling

= 1.47*10^9 g / 11821.4 g/s

=1.24*10^5 s

= 1.24*10^5 / 3600 hr

= 34.5 hr

Answer: 34.5 hr